Calculus, fall 2013

Study Guide for Chap 06

1. Practice on how to find A(x) in the formula ∫_{a}^{b} A(x)dx. Read the examples in and section 6.1 and try a few in Exercise 6.1. Also think about how to choose the best direction of slicing in these examples.

2. Study the method of disks and method of cylindrical shells for both volume of revolution around x− and y− axis.

3. Study the formula for arclength and area of surface of revolution around x− and y− axis.